OpenStudy (anonymous):

What is the value of y? 6(5 + 3) = 6 • 5 + 6 • y y =

jhonyy9 (jhonyy9):

30+18=30+6y 18=6y y=18/6 y=3

OpenStudy (anonymous):

you can't give full answers

jhonyy9 (jhonyy9):

this is the full answer